# Deep Learning Calabi-Yau four folds with hybrid and recurrent neural network architectures

### v2

@@ 6 Training results: 5-fold cross validation
In this section, we report the results obtained from training a single model, LSTM-448, using 5-fold cross validation (CV) on the combined training and validation sets from the 72% dataset. For the cross-validation experiments, we initially chose two models: CNN-LSTM-400 and LSTM-448, but CNN-LSTM-400 training was terminated early due to its slow convergence on the first fold of data. The training process is the same as that used for training the 12 models of Table 1 and has been described in detail in the previous section. In each fold, a new LSTM-448 model was trained from scratch using the dataset corresponding to that fold (see the discussion on the data preparation in Section 2). The final outcomes of the 5-fold CV experiment are the five different LSTM-448 models LSTM-448-f0, LSTM-448-f1, LSTM-448-f2, LSTM-448-f3, and LSTM-448-f4.
@@ 6.2 Ensembles
Compared to the ensembles of Table 5 formed using the models trained on the 72% dataset, the ensembles of Table 7 provide a significant improvement in the results. This can be visualized in Fig.23 in which the accuracies of the models in Table 7 and Table 6 are ranked together with the top four performing models from the first round of training (Ens-2, Ens-3, Ens-1 from Table 5, CNN-LSTM-400 from Table 3). In terms of the accuracies as listed in Table 7 above, the best ensemble is Ens-0f1f4-CL400, whose accuracy performance exceeds the best ensemble Ens-2 created from the models trained on the 72% dataset. Additional model rankings and performances can be found in Table 17 and Fig.30 in terms of the MSE metric, Table 18 and Fig. 31 in terms of the MAE metric, Table 19 and Fig.32 in terms of the $R^{2}$ metric, Table 20 and Fig.33 for all three metrics.
@@ 7.1 Individual models
Figure 24: Validation losses of the three models in normal and log scales.
@@ 8 Summary & Outlook
In this work, we explored and experimented with hybrid convolutional-recurrent and purely recurrent neural network architectures for the task of deep learning the Hodge numbers of CICY4 dataset. After carrying out multiple experiments involving training twelve models listed in Table 1 at 72% training ratio (including the 5-fold cross validation training of LSTM-448 in Table 6) and retraining the top three models listed in Table 8 at 80% training ratio, as well as forming various ensembles of the best performing models (Tables 5, 7, 9), our main findings are summarized below.
@@
  • For the hybrid CNN-RNN models, the CNN block, containing only about 10% of the total trainable parameters, is designed to be simple and light weight compared to the RNN block, which is intended to play the more dominant role. Two variants of CNN block were used: a custom CNN network with two convolutional layers, and a ResNet-inspired CNN network with two branches, a main branch containing three convolutional layers, and a short-cut branch with a single dense layer (with no bias). Due to the longer training time of hybrid networks utilizing the ResNet block, the majority of the CNN-RNN hybrid models use the custom CNN block. Varying the hidden size $M$ of the RNN block with the RNN being either GRU or LSTM let us observe the effect of $M$ on the model performance. The general trend observed is an increase in the obtained accuracy of the Hodge number predictions when $M$ is increased. When the peak performance is achieved at a certain $M$, increasing $M$ further leads to a decrease in performance or a failure to converge. Furthermore, it is observed that CNN-LSTM models significantly outperform CNN-GRU models, a result perhaps attributable to the fact that LSTMs in general often outperform GRUs. The results of CNN-RNN hybrid models can be found in Tables 2, 3.
@@
  • For the purely recurrent LSTM-based models, their observed performances are comparable to those CNN-LSTM models of roughly the same size (see Table 4). Similar to the trend observed with CNN-RNN, increasing the size of $M$ leads to an increase in performance until a peak is reached, after which performance drops.
@@
  • In comparison with the work [25] in which the authors employed a specialized inception-based network with around $10^{7}$ parameters (whose training time took approximately 5 days on a single NVIDIA V100 GPU) and obtained the results of 100% for both $h^{(1,1)}$ and $h^{(2,1)}$, 96% for $h^{(3,1)}$ and 83% for $h^{(2,2)}$ at 80% training ratio, our results are really promising, given the fact that our best models (LSTM-448, :ST-448-f1, LSTM-424, CNN-LSTM-400) are all around three times smaller and requiring much shorter training time (around 7 hours fully on the Kaggle P100 GPU for both rounds of training). While we did not achieve 100% accuracy for either $h^{1,1}$ or $h^{2,1}$, we did achieve better accuracies for $h^{3,1}$ and $h^{2,2}$ compared to [25] for both the case of an individual model (LSTM-448 at 80% training ratio) and the case of ensembles (Ens-0f1f4-CL400 at 72% and Ens-80-4 at 80% training ratios). This demonstrates the suitability and effectiveness of recurrent-based models (both in the pure and hybrid forms) for this specific task of regressing the Hodge numbers.
@@
Finally, we would like to conclude this work by noting that given the relatively simple choice of our neural network architectures, there is still a lot of room for experimentation and improvement. The reported results in this note are by no means the best possible ones that could be obtained by these model architectures. In particular, since the CNN part of the hybrid models is light weight and simple, its structure could be modified (either by adding more convolutional layers or changing the numbers of filters in the existing convolutional layers) and tuned, while keeping the RNN structure fixed, to observe the change in the results. The same kind of modifications can be applied to the ResNet-inspired variant of the CNN-RNN hybrid models. For the RNN block, we have kept the numbers of layers $L$ fixed to be 2 throughout this work, but it might be interesting to vary this number to observe its impact. During our initial experimentations, we observed that choosing $L=2$ yieled better results than $L=1$ or $L=3$ (or higher), but this might change when a different CNN structure is considered. For both the hybrid and purely recurrent architectures, there is an option to make the RNNs bidirectional but we decided not to, because bidirectionality will double the number of parameters in the network and increase the training time significantly. Furthermore, due to the constraint of the GPU usage allowed by the online platform Kaggle, the systematic scanning (parameter optimization by a dedicated library such as Optuna [^12] https://optuna.org/) [54] of the hidden size $M$ of all GRU/LSTM networks to find the optimum $M$ at which the highest accuracy can be obtained could not be carried out. [^13] To be able to perform an effective scan for $M$, a sizeable portion (in the order of at least tens of thousands of samples) of the training set should be used, which leads to very long scanning time well exceeding the allowed GPU time limit. It would be useful to perform a scanning for the RNN hidden size $M$ (and perhaps the number of layers $L$) if there were a better access to GPU resources.
@@
A more drastic change would involve doing away with convolutional or recurrent structures altogether in favor of experimenting with the transformer-based architecture [46], [47] which has been rapidly replacing both CNNs in computer vision and RNNs in natural language processing as the state-of-the-art technique. An approach that we would take when using a transformer-based architecture for this task of learning the Hodge numbers involves treating the CICY4 configuration matrices as images and decomposing them into patches that can be fed into a vision transformer (ViT) neural network [47] [^14] The ViT network is a custom network written from scratch using PyTorch, not a pretrained one. An experiment implementing this is already under way, but the initial indications are not promising - a single epoch of training using ViT takes about five minutes, about five times longer than the longest time taken by the largest model (CNN-LSTM-416)151515using the same Kaggle P100 GPU. Moreover, given the fact that due to the lack of certain inductive biases (such as translation equivariance and locality), transformer-based networks do not perform as well as CNNs when the dataset contains less than one million data points [47] (which is the case with the CICY4 dataset), we might already have an early indication of the result of the experiment involving transformer-based models.
@@ Acknowledgements
This work is possible thanks to the publicly free and available GPU cloud computing resources offered by Kaggle (https://www.kaggle.com).
@@ 9.3 MSE, MAE, R-squared metrics (80% dataset)
9.4 Training curves for all neural networks
@@
Figure 42: Train and validation losses of all models in Table 8 trained on the enlarged 80% dataset. The models were trained with the starting point being the saved checkpoints of the same models after being trained on the 72% dataset.

### v3

@@ 6 Training results: 5-fold cross validation
In this section, we report the results obtained from training a single model, LSTM-448, using 5-fold cross validation (CV) on the combined training and validation sets from the 72% dataset. For the cross-validation experiments, we initially chose two models: CNN-LSTM-400 and LSTM-448, but CNN-LSTM-400 training was terminated early due to its slow convergence on the first fold of data. The training process is the same as that used for training the 12 models of Table 1 and has been described in detail in the previous section. In each fold, a new LSTM-448 model was trained from scratch using the dataset corresponding to that fold (see the discussion on the data preparation in Section 2). For the distributions of the training and validation data in each fold, see Appendix 9.4. The final outcomes of the 5-fold CV experiment are the five different LSTM-448 models LSTM-448-f0, LSTM-448-f1, LSTM-448-f2, LSTM-448-f3, and LSTM-448-f4.
@@ 6.2 Ensembles
Some further comments are in order regarding the ensembles formed above:
  • While it is a common practice to include the results of all folds in a cross validation training to form an ensemble, a well proven technique in machine learning competitions (as hosted on Kaggle, see for example the discussion in [56]) is to form weighted ensembles where the models trained on different CV folds are weighted differently to obtain the best results when evaluated on the test dataset. There are various ways to obtain the optimal weights to be used in the weighted ensemble [55]. A very popular method is ‘hill climbing’ that starts with the best model and cycles through the list of all other available models together with a list of weights between a certain range to pick the second best model with a certain weight to form the first ensemble. Next, the third best model with a certain weight is chosen by repeating the same process of cycling through the remaining models to form a new ensemble with the models chosen from the previous iterations. The process continues until the results stop improving, i.e. added models no longer change the obtained results.
  • In the described process above, models resulting from CV training can be treated independently [^14] In a practical sense, a neural network model is nothing but a collection of its weights and biases so a different set of weights and biases constitutes an entirely different model. Models trained 5-fold CV are in fact independent models following this practical sense. and added to the mix in the same way as a model not trained using CV, as long as all models are evaluated on the same test set. The result of this exercise might lead to some optimized ensembles with certain CV folds not included in the final mix. In this work, due to the time and computational resource constraints, we did not carry out the optimization of the weights to form the most optimized weighted ensembles. Instead, we heuristically form the ensembles above by selecting the top performing models from both CV and non-CV training.
@@
Compared to the ensembles of Table 5 formed using the models trained on the 72% dataset, the ensembles of Table 7 provide a significant improvement in the results. This can be visualized in Fig.23 in which the accuracies of the models in Table 7 and Table 6 are ranked together with the top four performing models from the first round of training (Ens-2, Ens-3, Ens-1 from Table 5, CNN-LSTM-400 from Table 3). In terms of the accuracies as listed in Table 7 above, the best ensemble is Ens-0f1f4-CL400, whose accuracy performance exceeds the best ensemble Ens-2 created from the models trained on the 72% dataset. The second best model is Ens-f1f4-CL400 whose accuracy also exceeds Ens-2. The difference between Ens-0f1f4-CL400 and Ens-f1f4-CL400 is the addition of the non-CV model LSTM-448. While it might appears counterintuitive that the addition of LSTM-448, whose performance ranks below CNN-LSTM-400 on the 72% dataset, elevates the performance of Ens-0f1f4-CL400 above Ens-f1f4-CL400, it is perfectly logical once one recalls that different models have different regions of competence that might be complimentary to one another’s. This is why ensembles often outperform their constituent models due to their robustness and better capability to generalize when evaluated on the same test set.  [^15] Heuristically, it is not impossible to have an ensemble (let’s call it Ens-AB) that is made of Model A with 55% accuracy and Model B with 65% accuracy obtaining 70% accuracy. It is again not impossible that adding another model, Model C with 60% accuracy, leads to an even better ensemble (let’s call it Ens-ABC) with 72% accuracy. In this case, Model B at 65% accuracy clearly outperforms Model C (at 60%), but the new ensemble Ens-ABC including Model C can outperforms the old ensemble (Ens-AB). In AI/ML competitions, this is a scenario that happens repeatedly and winning solutions often make use of ensembles of multiple models in this exact sense ([57], [58], [59]). Additional model rankings and performances can be found in Table 17 and Fig.30 in terms of the MSE metric, Table 18 and Fig. 31 in terms of the MAE metric, Table 19 and Fig.32 in terms of the $R^{2}$ metric, Table 20 and Fig.33 for all three metrics.
@@ 7.1 Individual models
Figure 24: Validation losses (in log scale) of the three models trained from scratch listed in Table 8.
@@ 8 Summary & Outlook
In this work, we explored and experimented with hybrid convolutional-recurrent and purely recurrent neural network architectures for the task of deep learning the Hodge numbers of CICY4 dataset. We emphasize that our work is by nature a proof-of-concept type, exploring and establishing the utility of RNN-based architectures for the task of regressing or learning the Hodge numbers, rather than a full-fledged characterization of which RNN architectures perform best. In this sense, it should be noted that while we used [25] as a benchmark to compare all our RNN-based models against, there are significant differences in the way the problem was approached. Unlike [25], we did not start with four separate neural network models to learn each Hodge number independently161616as done in [25], the goal of this was to find the best architecture and parameters on which to base the final model, ‘CICYMiner’, capable of learning all 4 Hodge numbers simultaneously, but chose to begin instead with a single RNN-based model (with a different architecture variant in each experiment) to learn all four Hodge numbers simultaneously. This approach was adopted based on the context that CNN-based architectures were already the established and default type of neural networks to use for this type of problem since 2017 ([2]), and it was only a matter of finding out which CNN architecture was the best to use in subsequent works. In this work, the starting point of the investigation was more fundamental, in the sense that RNN-based architectures had never been applied to this kind of problem, so the natural question that arose was: Can we apply an RNN to this and will we obtain similar results to those obtained by CNNs? In this context, the most computationally efficient way to carry out this task is to create one single deep-learning model to regress all four Hodge numbers, and contrast the performance of this model with the best-performing CNN model already established. After carrying out multiple experiments involving training twelve models listed in Table 1 at 72% training ratio (including the 5-fold cross validation training of LSTM-448 in Table 6) and retraining the top three models listed in Table 8 at 80% training ratio, as well as forming various ensembles of the best performing models (Tables 5, 7, 9), our main findings are summarized below.
@@
  • For the hybrid CNN-RNN models, the CNN block, containing only about 10% of the total trainable parameters, is designed to be simple and light weight compared to the RNN block, which is intended to play the more dominant role. Two variants of CNN block were used: a custom CNN network with two convolutional layers, and a ResNet-inspired CNN network with two branches, a main branch containing three convolutional layers, and a short-cut branch with a single dense layer (with no bias). Due to the longer training time of hybrid networks utilizing the ResNet block, the majority of the CNN-RNN hybrid models use the custom CNN block. Varying the hidden size $M$ of the RNN block with the RNN being either GRU or LSTM let us observe the effect of $M$ on the model performance. The general trend observed is an increase in the obtained accuracy of the Hodge number predictions when $M$ is increased. When the peak performance is achieved at a certain $M$, increasing $M$ further leads to a decrease in performance or a failure to converge, the cause of which is likely to be overfitting. Furthermore, it is observed that CNN-LSTM models significantly outperform CNN-GRU models, a result perhaps attributable to the fact that LSTMs in general often outperform GRUs due to their more sophisticated recurrent architecture. The results of CNN-RNN hybrid models can be found in Tables 2, 3.
@@
  • For the purely recurrent LSTM-based models, their observed performances are comparable to those CNN-LSTM models of roughly the same size (see Table 4). Similar to the trend observed with CNN-RNN, increasing the size of $M$ leads to an increase in performance until a peak is reached, after which performance drops due to overfitting.
@@
  • In comparison with the work [25] in which the authors employed a specialized inception-based network with around $10^{7}$ parameters (whose training time took approximately 5 days on a single NVIDIA V100 GPU) and obtained the results of 100% for both $h^{(1,1)}$ and $h^{(2,1)}$, 96% for $h^{(3,1)}$ and 83% for $h^{(2,2)}$ at 80% training ratio, our results are really promising, given the fact that our best models (LSTM-448, LSTM-448-f1, LSTM-424, CNN-LSTM-400) are all around three times smaller and requiring much shorter training time (around 7 hours fully on the Kaggle P100 GPU for both rounds of training). While we did not achieve 100% accuracy for either $h^{1,1}$ or $h^{2,1}$, we did achieve better accuracies for $h^{3,1}$ and $h^{2,2}$ compared to [25] for both the case of an individual model (LSTM-448 at 80% training ratio) and the case of ensembles (Ens-0f1f4-CL400 at 72% and Ens-80-4 at 80% training ratios). This demonstrates the suitability and effectiveness of recurrent-based models (both in the pure and hybrid forms) for this specific task of regressing the Hodge numbers.
@@
Finally, we would like to conclude this work by noting that given the relatively simple choice of our neural network architectures, there is still a lot of room for experimentation and improvement. The reported results in this paper are by no means the best possible ones that could be obtained by these model architectures. In particular, since the CNN part of the hybrid models is light weight and simple, its structure could be modified (either by adding more convolutional layers or changing the numbers of filters in the existing convolutional layers) and tuned, while keeping the RNN structure fixed, to observe the change in the results. The same kind of modifications can be applied to the ResNet-inspired variant of the CNN-RNN hybrid models. For the RNN block, we have kept the numbers of layers $L$ fixed to be 2 throughout this work, but it might be interesting to vary this number to observe its impact. During our initial experimentations, we observed that choosing $L=2$ yieled better results than $L=1$ or $L=3$ (or higher), but this might change when a different CNN structure is considered. For both the hybrid and purely recurrent architectures, there is an option to make the RNNs bidirectional but we decided not to, because bidirectionality will double the number of parameters in the network and increase the training time significantly. Furthermore, due to the constraint of the GPU usage allowed by the online platform Kaggle, the systematic scanning (parameter optimization by a dedicated library such as Optuna [^17] https://optuna.org/) [54] of the hidden size $M$ of all GRU/LSTM networks to find the optimum $M$ at which the highest accuracy can be obtained could not be carried out. [^18] To be able to perform an effective scan for $M$, a sizeable portion (in the order of at least tens of thousands of samples) of the training set should be used, which leads to very long scanning time well exceeding the allowed GPU time limit. It would be useful to perform a scanning for the RNN hidden size $M$ (and perhaps the number of layers $L$) if there were a better access to GPU resources.
@@
A more drastic change would involve doing away with convolutional or recurrent structures altogether in favor of experimenting with the transformer-based architecture [46], [47] which has been rapidly replacing both CNNs in computer vision and RNNs in natural language processing as the state-of-the-art technique. An approach that we would take when using a transformer-based architecture for this task of learning the Hodge numbers involves treating the CICY4 configuration matrices as images and decomposing them into patches that can be fed into a vision transformer (ViT) neural network [47] [^19] The ViT network is a custom network written from scratch using PyTorch, not a pretrained one. An experiment implementing this is already under way, but the initial indications are not promising - a single epoch of training using ViT takes about five minutes, about five times longer than the longest time taken by the largest model (CNN-LSTM-416)202020using the same Kaggle P100 GPU. Moreover, given the fact that due to the lack of certain inductive biases (such as translation equivariance and locality), transformer-based networks do not perform as well as CNNs when the dataset contains less than one million data points [47] (which is the case with the CICY4 dataset), we might already have an early indication of the result of the experiment involving transformer-based models.
@@ Acknowledgements
The author is an unaffiliated, independent researcher - this work was possible thanks solely to the publicly free and available GPU cloud computing resources offered by Kaggle (https://www.kaggle.com).
@@ 9.3 MSE, MAE, R-squared metrics (80% dataset)
9.4 Distributions of the Hodge numbers in 5-fold cross validation training
Figure 38: Distributions of the Hodge numbers in the first fold of 5-fold CV training
Figure 39: Distributions of the Hodge numbers in the second fold of 5-fold CV training
Figure 40: Distributions of the Hodge numbers in the third fold of 5-fold CV training
Figure 41: Distributions of the Hodge numbers in the fourth fold of 5-fold CV training
Figure 42: Distributions of the Hodge numbers in the fifth fold of 5-fold CV training
9.5 Training curves of all neural networks
@@
Figure 47: Train and validation losses of the 3 models in Table 8 trained from scratch on the enlarged 80% dataset. The models were trained with the starting point being the saved checkpoints of the same models after being trained on the 72% dataset.